# {refsection}

### v2

@@
Yuri Z. Sinzato111y.z.sinzato@uva.nl, ORCID: $0$-0002-3324-5712 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Robert Uittenbogaard Hydro-Key Ltd, Haelen 6081EA, The Netherlands Petra M. Visser Department of Freshwater and Marine Ecology, Institute for Biodiversity and Ecosystem Dynamics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Jef Huisman Department of Freshwater and Marine Ecology, Institute for Biodiversity and Ecosystem Dynamics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Maziyar Jalaal222m.jalaal@uva.nl, ORCID: $0$-0002-5654-8505 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands
@@
Keywords: Algal blooms |||| Cyanobacteria |||| Hydrodynamics |||| Microcystis
@@ Introduction
Microcystis cells are held together by a layer of extracellular polymeric substances (EPS), facilitating the formation of large colonies of up to a millimeter in diameter with diverse morphological arrangements [van2022microcystis, liu2018characteristics, via2004distribution]. Two main mechanisms for colony formation have been proposed [xiao2018colony]: (i) cell division, i.e., clonal expansion of cells which remain attached by their secreted EPS layer after division, and (ii) aggregation, i.e., initially independent cells/colonies collide and adhere to each other. Sequencing of Microcystis colonies has revealed a much higher genetic diversity between independent colonies than within colonies, thus supporting the hypothesis that colony formation under natural conditions is primarily driven by cell division [perez2021single]. However, cell aggregation is a possible route for increase in colony size on a short time scale, and may offer improved protection against high grazing pressure [xiao2017meta]. Aggregation becomes significant when the collision frequency is high enough for aggregation rates to overcome cell division rate [burd2009particle], and the stickiness of colonies is sufficiently large (controlled by the EPS composition, pH, and ionic strength of water or medium [van1988aggregation, kiorboe1993phytoplankton, verspagen2006aggregation]). Yet, it remains an open question to what extent aggregation of Microcystis cells and colonies is a relevant process in natural systems. In addition to aggregation, fluid flow also influences colony formation through fragmentation, potentially serving as a mechanism for colony reproduction [day2022varied] to overcome the growth limitations of large colonies [feng2020structural]. Beyond modulating colony size through aggregation and fragmentation, fluid flow has also been proposed as a driving factor in changes in colony morphology and growth rate [li2018morphospecies, wilkinson2016effect, xiao2016effect].
@@
The results show that the initial size distribution had a bimodal shape, with a subpopulation of small colonies peaking around $l\sim 1-2$ (single cells, dimers, and trimers) and a subpopulation of large colonies, peaking around $l\sim 20-40$ (Fig. 2A). The cut-off size between the two subpopulations is defined here as the minimum of the bimodal distribution. Some small colonies ($l\sim 1-2$), up to $20\%$ of the total biovolume, were always present in the initial distribution even after filtration, due to pre-stirring required to keep the suspension homogeneous. However, medium-sized colonies ($l=4-8$) accounted for a negligible fraction of the distribution.
@@
Direct measurement of the fragment distribution function $g(l,l^{\prime})$ (i.e., the frequency of fragments with size $l$ arising from the fragmentation of a colony of size $l^{\prime}$) is an experimental challenge [flesch1999laminar, vankova2007emulsification, haakansson2020experimental]. Individual fragmentation events could not be visualized as the residence time of colonies inside the field of view is too short. Instead, a suitable model can be identified from the rate of change in the biovolume distribution (Fig. 2E). A negative rate was observed for large colonies in the range $l\gtrsim 25$, indicating a loss in that size range by fragmentation. In contrast, the newly created fragments appeared as positive peaks, with a distinct peak close to the single-cell size. This result is compatible with an erosion mechanism, in which small fragments with one or a few cells are detached from the outer layer of the colony. For an ideal binary erosion, a fragmentation event of a colony with $i$ cells results in one cell being eroded from the surface of the colony with a complementary fragment of $i-1$ cells remaining. Small deviations from an ideal binary erosion are observed in Figure 2E, with most eroded fragments lying in the range $1\leq l\leq 2$. Nevertheless, ideal erosion provides a good approximation for the results. Accurate identification of complementary fragments from the rate of change in the biovolume distribution was more difficult, because the broad initial distribution of our suspension led to a broad distribution of complementary fragments. In this case, the rate of change in the biovolume distribution shifted from a net loss of eroded colonies at larger sizes ( $l\gtrsim 25$) to a net gain of complementary fragments at intermediate sizes ( $5\lesssim l\lesssim 20$).
@@
Fig. 2: Kinetics of fragmentation of Microcystis strain V163 colonies under cone-and-plate shear flow. The laboratory culture was filtered to select mainly large colonies, and the total biovolume fraction was adjusted to $\phi=10^{-4}$. Suspensions were subjected to an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$) in panels A-E. (A) The initial size distribution of colonies, expressed as biovolume fraction of the relative colony diameter (normalized by single cell diameter). The size distribution had a bimodal shape, with large colonies (in yellow) and small colonies (in green, composed mostly of single cells, dimers, and some trimers, as depicted by the inset). (B) Median diameter of small and large colonies as a function of time. Symbols indicate the experimental data, while the lines indicate the predictions from the population model given by Eq. (1) with an erosion (dashed) or equal fragments (dot-dash) hypothesis for large division-formed colonies. Small aggregates colonies follow the equal fragments hypothesis. Bars and shaded region indicate limits of 25th and 75th percentiles. (C) Most large colonies have been fragmented after 1 hour of shear flow, but the bimodal shape of the colony size distribution remained. (D) Biovolume fraction of small colonies (i.e., biovolume of small colonies over the total biovolume) as a function of time. A shift is observed from large to small colonies, captured well by the erosion model, but not by the equal fragments model. (E) The rate of change in biovolume distribution at $t=0$h. Negative values indicate loss of colonies by fragmentation, while positive values indicate newly created fragments. The distribution suggests an erosion mechanism, as depicted by the cartoon inside the plot. (F) Fragmentation frequency as a function of the relative diameter of colonies for three values of dissipation rate. The inset shows details for moderate dissipation rates. Error bars indicate the standard deviation and the lines indicate the predictions by Eq. (5). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$; erosion: $S_{2}=31$ , $q_{2}=4.1$ ; equal fragments: $S_{2}=33$ , $q_{2}=6.3$.
@@
The results above indicate that bonds formed during flow-induced aggregation (whether between two single cells or two division-formed colonies) are weak and cannot withstand intense dissipation rates. In contrast, the bonds between cells within a single division-formed colony are significantly more resistant. This difference in mechanical strength of the cellular binding in aggregation-formed versus division-formed colonies can be attributed to the structure of the EPS layer that surrounds the colonies. Aggregated colonies would have intercellular gaps (smaller that the cell size) that are not filled with EPS. In contrast, cell division allows enough time for the gaps to be filled with secreted EPS, effectively increasing the bond strength between cells (Figs. 3B and 3D). However, we could not confirm this hypothesis visually, as the EPS layer in Microcystis strain V163 is too thin ($<1\mu m$ ) for a clear identification using optical microscopy.
@@
We solved the balance equation (1) using a zero-order discrete population balance, or sectional method, in which the size distribution is described by uniformly sized bins where the size distribution is evaluated at a representative size centered on the bin [vanni2000approximate, kumar1996solution]. The simulations were initialized with the size distributions at $t=0$ obtained in our experiments. The two pairs of parameters for critical size ({$q_{1}$,$S_{1}$} and {$q_{2}$,$S_{2}$}) and the stickiness parameter ($\alpha_{1}$) were used to fit the model to the experimental data (Materials and Methods).
@@
For an initial suspension composed mainly of large colonies at an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$), the model prediction of a decrease in the median size of $C_{2}$ colonies up to a steady-state diameter is in good agreement with the experimental findings (Fig. 2B). Moreover, the erosion mechanism accurately predicts the biovolume transfer from large to small colonies (Fig. 2D). In comparison, the hypothesis of equal fragments formation for $C_{2}$ colonies does not capture such transfer, as all fragments remain in the large colony range (Fig. 2D and Fig. S9 of SI Appendix).The use of the erosion mechanism for the fragment distribution function is further corroborated by previous studies simulating bacterial aggregates [byrne2011postfragmentation]. For $C_{1}$ colonies at this intense dissipation rate, the model predicts that their median diameter remains near $l^{*}=1$ (i.e., single cells). However, the experimental results show a slightly wider size distribution of both the small and large colonies than predicted by an ideal erosion mechanism (Fig. 2C). Nevertheless, the global behavior of the size distribution is well described by the two-category model. Furthermore, a very good agreement is also observed between the measured fragmentation frequency and the model predictions given by Eq. (5) (Fig. 2F).
@@
To complement the mathematical model and to further investigate the mechanical origins of the resistance of division-formed colonies to intense shear, we performed rheology tests on the cyanobacterial colonies (Materials and Methods). Simple and oscillatory shear measurements revealed elasto-viscoplastic mechanics [balmforth2014yielding, bonn2017yield] of the concentrated aggregate with a yield stress of $\tau_{y}=4.3\pm 0.3\text{ Pa}$. The low volume fraction of cells in aggregate ($\sim 12$) indicates that the mechanical properties are mainly dictated by the EPS layer. Equating the average hydrodynamic shear stress to the yield stress in Eq. (11) and subsequently substituting the critical dissipation rate in Eq. (3) results in a critical colony size of $l^{*}_{2}=2.8$. The eroded fragments observed in the fragmentation experiments (Fig. 2E) are mostly smaller than this critical size. This suggests that large colonies will suffer erosion when the local hydrodynamic stress exceeds the yield stress of the EPS layer. Although the fracture behavior of colonies in suspension is expected to be different from a concentrated aggregate under shear rheology, the latter provides a reasonable and convenient estimate of the material properties of the colonies. Related to the critical size of the colony and their mechanical properties is the exponent $q_{2}$ (see equation 5). The best fit of the results in Fig. 2 provided $q_{2}=4.1$. This value lies in the upper limit of previous results for colloidal aggregates [flesch1999laminar, baebler2008modelling]. Note that for colloids, $q$ is related to the stress concentration around cracks in the fractal structure of the aggregate [zaccone2009breakup], and low values of $q$ are associated with brittle deformation, in which the stress concentration grows quickly with aggregate size. In contrast, cyanobacterial colonies have a soft plastic behavior that is less sensitive to large cracks, thus justifying the large value of $q$.
@@
[…]
@@ Conclusions
We have demonstrated that flow-induced aggregated colonies of Microcystis form weaker structures than division-formed colonies, likely due to the weak EPS bonds between cells. Furthermore, we showed that colony aggregation is unlikely in lakes under early bloom conditions, as the collision frequency of colonies is too low. Our results, based on the mechanical properties of colonies, are corroborated by previous studies based on the sequencing of individual Microcystis colonies [perez2021single, smith2021individual]. These studies have identified that the genetic diversity within Microcystis colonies is much smaller than the diversity between Microcystis colonies, thus supporting the hypothesis that colony formation is dominated by cell division rather than by aggregation of colonies. Nonetheless, our model supports recent studies which demonstrated that, during scum formation in very dense blooms or phytoplankton patches, flow-induced aggregation likely plays a role and may quickly increase colony size [wu2019effects, wu2020recovery]. A summary of these outcomes is presented in the phase map illustrated in Figure 5.
@@ B.   Cone-and-plate shear
Our setup consists of an upper conical probe made from transparent acrylic, with an outer diameter of 42 mm and a cone complementary angle of $\theta=5.44^{o}$ (Fig. S7A). The conical probe is attached to a steel shaft, and it is driven by a rheometer head Anton Paar DSR 502. The lower surface consists of a 1mm thick glass slide. A cylindrical acrylic ring confines the fluid, preventing spillage of the sample due to strong centrifugal forces. The inner diameter of the ring is 46 mm, which ensures that the nominal shear rate between the upper probe’s outer surface and the ring is the same as between the conical and lower surfaces. A clearance gap of $10~{}\mu$m was set between the tip of the cone and the lower surface to avoid solid-solid friction.
@@
[…]

### v3

# Fragmentation and aggregation of cyanobacterial colonies

@@
Yuri Z. Sinzato111y.z.sinzato@uva.nl, ORCID: 0000-0002-3324-5712 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Robert Uittenbogaard Hydro-Key Ltd, Haelen 6081EA, The Netherlands Petra M. Visser Department of Freshwater and Marine Ecology, Institute for Biodiversity and Ecosystem Dynamics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Jef Huisman Department of Freshwater and Marine Ecology, Institute for Biodiversity and Ecosystem Dynamics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands Maziyar Jalaal222m.jalaal@uva.nl, ORCID: 0000-0002-5654-8505 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Science Park 904, Amsterdam, 1098XH, The Netherlands
@@
Keywords: Algal blooms $|$ Cyanobacteria $|$ Hydrodynamics $|$ Microcystis
@@ Introduction
Microcystis cells are held together by a layer of extracellular polymeric substances (EPS), facilitating the formation of large colonies of up to a millimeter in diameter with diverse morphological arrangements [van2022microcystis, liu2018characteristics, via2004distribution]. Two main mechanisms for colony formation have been proposed [xiao2018colony]: (i) cell division, i.e., clonal expansion of cells which remain attached by their secreted EPS layer after division, and (ii) aggregation, i.e., initially independent cells/colonies collide and adhere to each other. Sequencing of Microcystis colonies has revealed a much higher genetic diversity between independent colonies than within colonies, thus supporting the hypothesis that colony formation under natural conditions is primarily driven by cell division [perez2021single]. In contrast, cell aggregation (sometimes also called cell adhesion) can promote a rapid increase in colony size beyond the limit set by division rates, and may explain sudden rises in colony size in late bloom periods [xiao2017meta, zhu2015differences, li2013seasonal]. Aggregation rates depend on the stickiness of the colonies, which in turn is controlled by the EPS composition, pH, and ionic composition of water [van1988aggregation, kiorboe1993phytoplankton, verspagen2006aggregation]. In particular, divalent cations such as Ca2+ can bridge negatively charged functional groups in EPS and therefore increase stickiness [xu2016electrolyte, wang2011effects, sato2017colony]. It has been shown that high levels of Ca2+ enhance cell aggregation in Microcystis cultures [chen2020calcium]. Moreover, cell aggregation can provide a fast defense against grazing [yang2006morphological]. Fluid flow plays an important role in cell aggregation by regulating the collision frequency between cells or colonies [burd2009particle]. In addition, fluid flow also influences colony formation through fragmentation, potentially serving as a mechanism for colony reproduction [day2022varied] to overcome the growth limitations of large colonies [feng2020structural]. Beyond modulating colony size through aggregation and fragmentation, fluid flow has also been proposed as a driving factor in changes in colony morphology and growth rate [li2018morphospecies, wilkinson2016effect, xiao2016effect].
@@
The results show that the initial size distribution had a bimodal shape, with a subpopulation of small colonies peaking around $l\sim 1-2$ (single cells, dimers, and trimers) and a subpopulation of large colonies, peaking around $l\sim 20-40$ (Fig. 2A). The cut-off size between the two subpopulations is defined here as the minimum of the bimodal distribution. Some small colonies ($l\sim 1-2$), up to $20\%$ of the total biovolume, were always present in the initial distribution even after filtration, due to pre-stirring required to keep the suspension homogeneous. However, medium-sized colonies (l=48l=4-8) accounted for a negligible fraction of the distribution.
@@
Direct measurement of the fragment distribution function $g(l,l^{\prime})$ (i.e., the frequency of fragments with size $l$ arising from the fragmentation of a colony of size $l^{\prime}$) is an experimental challenge [flesch1999laminar, vankova2007emulsification, haakansson2020experimental]. Individual fragmentation events could not be visualized as the residence time of colonies inside the field of view is too short. Instead, a suitable model can be identified from the rate of change in the biovolume distribution (Fig. 2E). A negative rate was observed for large colonies in the range $l\gtrsim 25$, indicating a loss in that size range by fragmentation. In contrast, the newly created fragments appeared as positive peaks, with a distinct peak close to the single-cell size. This result is compatible with an erosion mechanism, in which small fragments with one or a few cells are detached from the outer layer of the colony. For an ideal binary erosion, a fragmentation event of a colony with $i$ cells results in one cell being eroded from the surface of the colony with a complementary fragment of i1i-1 cells remaining. Small deviations from an ideal binary erosion are observed in Figure 2E, with most eroded fragments lying in the range $1\leq l\leq 2$. Nevertheless, ideal erosion provides a good approximation for the results. Accurate identification of complementary fragments from the rate of change in the biovolume distribution was more difficult, because the broad initial distribution of our suspension led to a broad distribution of complementary fragments. In this case, the rate of change in the biovolume distribution shifted from a net loss of eroded colonies at larger sizes ( $l\gtrsim 25$) to a net gain of complementary fragments at intermediate sizes ( $5\lesssim l\lesssim 20$).
@@
Fig. 2: Kinetics of fragmentation of Microcystis strain V163 colonies under cone-and-plate shear flow. The laboratory culture was filtered to select mainly large colonies, and the total biovolume fraction was adjusted to $\phi=10^{-4}$. Suspensions were subjected to an intense dissipation rate ($\dot{\varepsilon}=5.8~{m^{2}/s^{3}}$) in panels A-E. (A) The initial size distribution of colonies, expressed as biovolume fraction of the relative colony diameter (normalized by single cell diameter). The size distribution had a bimodal shape, with large colonies ($l>5$, in yellow) and small colonies ($l\leq 5$, in green, composed mostly of single cells, dimers, and some trimers, as depicted by the inset). (B) Median diameter of small and large colonies as a function of time. Symbols indicate the experimental data, while the lines indicate the predictions from the population model given by Eq. (1) with an erosion (dashed) or equal fragments (dot-dash) hypothesis for large division-formed colonies. Small aggregates colonies follow the equal fragments hypothesis. Bars and shaded region indicate limits of 25th and 75th percentiles. (C) Most large colonies have been fragmented after 1 hour of shear flow, but the bimodal shape of the colony size distribution remained. (D) Biovolume fraction of small colonies (i.e., biovolume of small colonies over the total biovolume) as a function of time. A shift is observed from large to small colonies, captured well by the erosion model, but not by the equal fragments model. (E) The rate of change in biovolume distribution at t=0t=0h. Negative values indicate loss of colonies by fragmentation, while positive values indicate newly created fragments. The distribution suggests an erosion mechanism, as depicted by the cartoon inside the plot. (F) Fragmentation frequency as a function of the relative diameter of colonies for three values of dissipation rate. The inset shows details for moderate dissipation rates. Error bars indicate the standard deviation and the lines indicate the predictions by Eq. (5). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$; erosion: $S_{2}=31$ , $q_{2}=4.1$ ; equal fragments: $S_{2}=33$ , $q_{2}=6.3$.
@@
The results above indicate that bonds formed during flow-induced aggregation (whether between two single cells or two division-formed colonies) are weak and cannot withstand intense dissipation rates. In contrast, the bonds between cells within a single division-formed colony are significantly more resistant. This difference in mechanical strength of the cellular binding in aggregation-formed versus division-formed colonies can be attributed to the structure of the EPS layer that surrounds the colonies. Aggregated colonies would have intercellular gaps (smaller that the cell size) that are not filled with EPS. In contrast, cell division allows enough time for the gaps to be filled with secreted EPS, effectively increasing the bond strength between cells (Figs. 3B and 3D). However, we could not confirm this hypothesis visually, as the EPS layer in Microcystis strain V163 is too thin ($1\mu m$ ) for a clear identification using optical microscopy.
@@
We solved the balance equation (1) using a zero-order discrete population balance, or sectional method, in which the size distribution is described by uniformly sized bins where the size distribution is evaluated at a representative size centered on the bin [vanni2000approximate, kumar1996solution]. The simulations were initialized with the size distributions at $t=0$ obtained in our experiments. The two pairs of parameters for critical size ($q_{1}$$S_{1}$} and $q_{2}$$S_{2}$}) and the stickiness parameter ($\alpha_{1}$) were used to fit the model to the experimental data (Materials and Methods).
@@
For an initial suspension composed mainly of large colonies at an intense dissipation rate ($\dot{\varepsilon}=5.8~{m^{2}/s^{3}}$), the model prediction of a decrease in the median size of $C_{2}$ colonies up to a steady-state diameter is in good agreement with the experimental findings (Fig. 2B). Moreover, the erosion mechanism accurately predicts the biovolume transfer from large to small colonies (Fig. 2D). In comparison, the hypothesis of equal fragments formation for $C_{2}$ colonies predicts that all fragments remain within the large colony size category ($l>5$) during the first 5 hours of experiments, hence the equal fragments model does not capture the observed transfer from large to small colonies (Fig. 2D and Fig. S9 of SI Appendix).The use of the erosion mechanism for the fragment distribution function is further corroborated by previous studies simulating bacterial aggregates [byrne2011postfragmentation]. For $C_{1}$ colonies at this intense dissipation rate, the model predicts that their median diameter remains near $l^{*}=1$ (i.e., single cells). However, the experimental results show a slightly wider size distribution of both the small and large colonies than predicted by an ideal erosion mechanism (Fig. 2C). Nevertheless, the global behavior of the size distribution is well described by the two-category model. Furthermore, a very good agreement is also observed between the measured fragmentation frequency and the model predictions given by Eq. (5) (Fig. 2F).
@@
To complement the mathematical model and to further investigate the mechanical origins of the resistance of division-formed colonies to intense shear, we performed rheology tests on the cyanobacterial colonies (Materials and Methods). Simple and oscillatory shear measurements revealed elasto-viscoplastic mechanics [balmforth2014yielding, bonn2017yield] of the concentrated aggregate with a yield stress of $\tau_{y}=4.3\pm 0.3\text{ Pa}$. The low volume fraction of cells in aggregate ($\sim 12$ %) indicates that the mechanical properties are mainly dictated by the EPS layer. Equating the average hydrodynamic shear stress to the yield stress in Eq. (11) and subsequently substituting the critical dissipation rate in Eq. (3) results in a critical colony size of $l^{*}_{2}=2.8$. The eroded fragments observed in the fragmentation experiments (Fig. 2E) are mostly smaller than this critical size. This suggests that large colonies will suffer erosion when the local hydrodynamic stress exceeds the yield stress of the EPS layer. Although the fracture behavior of colonies in suspension is expected to be different from a concentrated aggregate under shear rheology, the latter provides a reasonable and convenient estimate of the material properties of the colonies. Related to the critical size of the colony and their mechanical properties is the exponent $q_{2}$ (see equation 5). The best fit of the results in Fig. 2 provided $q_{2}=4.1$. This value lies in the upper limit of previous results for colloidal aggregates [flesch1999laminar, baebler2008modelling]. Note that for colloids, $q$ is related to the stress concentration around cracks in the fractal structure of the aggregate [zaccone2009breakup], and low values of $q$ are associated with brittle deformation, in which the stress concentration grows quickly with aggregate size. In contrast, cyanobacterial colonies have a soft plastic behavior that is less sensitive to large cracks, thus justifying the large value of $q$.
@@
$\dfrac{1}{\overline{l}}\dfrac{\mathrm{d}\overline{l}}{\mathrm{d}t}=2.49(2^{1/f}-1)\,\phi\,\alpha\,\overline{l}^{\,4-2f}\sqrt{\dfrac{\dot{\varepsilon}}{\nu}}~.$ (8)
@@ Conclusions
We have demonstrated that flow-induced aggregated colonies of Microcystis form weaker structures than division-formed colonies, likely due to the weak EPS bonds between cells. A previous study on colony aggregation at high Ca2+ levels observed similar morphological differences in colony formation [chen2020calcium]. There, an initial fast cell aggregation produced a sparse colony structure, followed by a more compact structure of the colonies associated with cell division. Furthermore, we showed that colony aggregation is unlikely in lakes under early bloom conditions, as the collision frequency of colonies is too low. Our results, based on the mechanical properties of colonies, are corroborated by previous studies based on the sequencing of individual Microcystis colonies [perez2021single, smith2021individual]. These studies have identified that the genetic diversity within Microcystis colonies is much smaller than the diversity between Microcystis colonies, thus supporting the hypothesis that colony formation is dominated by cell division rather than by aggregation of colonies. Nonetheless, our model supports recent studies which demonstrated that, during scum formation in very dense blooms or phytoplankton patches, flow-induced aggregation likely plays a role and may quickly increase colony size [wu2019effects, wu2020recovery]. A summary of these outcomes is presented in the phase map illustrated in Figure 5.
@@ B.  Cone-and-plate shear
[…]
@@
Our cone-and-plate setup covers energy dissipation rates over three orders of magnitude, from $10^{-2}$ to $10~{m^{2}/s^{3}}$. At the lower end, the dissipation rate is limited by the sedimentation of colonies near the stagnation point at the conical tip. Due to the small gap, this stagnation point is not accessible to the colonies. A large colony of $l\sim 50$ does not fit within the surfaces up to a radial distance from the tip of $4$ mm. For the lowest dissipation rate ($\dot{\varepsilon}=0.019~{m^{2}/s^{3}}$), a minimum secondary flow speed of $\sim 100\mu\text{m/s}$ is present [sdougos1984secondary]. This value is above the sedimentation speed of colonies of Microcystis strain V163, which are either buoyant or slightly sinking due to the presence of gas vesicles. Furthermore, the region near the conical tip accounts for a small fraction of sample volume, while the velocity fluctuations in the bulk region of the cone-and-plate setup ensure that the colonies remain in suspension and are homogeneously distributed. The upper limit for the operating dissipation rate is set by the spillage of fluid out of the cylindrical confinement. Regarding the optimal range of biovolume fraction, this is limited at $\phi=0.0005$, above which the optical overlap between nearby colonies influences the colony count. The kinematic viscosity of the liquid $\nu$ can be taken equal to water for the biovolume fraction explored here.